# Self-Supervised Learning and Opportunistic Inference for Continuous Monitoring of Freezing of Gait in Parkinson’s Disease

### v1

@@ 4.2 Compare with State-of-the-Art Models
One Class Classifier [10] 90.8% 71.6% 0.86 0.77 0.89
Semi-Supervised Model [11] 72.3% 0.71
Multi-head CNN [12] 97.3% 94.6% 0.56 0.68 0.49
@@
4.3 Duration Based FoG Episode
@@
The SSL model shows robust performance across all durations, with detection rates increasing as the episode lengthens. This trend is mirrored in the baseline model’s performance, indicating a consistent improvement across different modeling approaches. For short episodes, the detection rate is 83.3% for episodes and 68% for windows, with an average latency of 2.42±0.45 seconds. For long episodes, these rates rise to 100% for episodes and 91.1% for windows, with a latency of 2.64$\pm$1.45 seconds. The increasing detection rates for longer episodes are due to the more prominent and persistent FoG characteristics, such as tremors and shuffling gait, which the SSL model effectively captures. The increasing detection rates for longer episodes are due to the more prominent and persistent FoG characteristics, such as tremors and shuffling gait, which LIFT-PD effectively captures.
Table 8: Duration-based FoG episode analysis.
Duration FoG Episodes & windows detection rate (%) with latency (s) and standard deviation
Small, 0–6 s Medium, 6–12 s Large, >12 s
FoG Episodes 83.3% (82.8%) 100% (98.8%) 100% (100%)
FoG Windows 68% (71.4%) 81.9% (84.1%) 91.1% (92.6%)
Avg. Latency $\pm$ SD 2.42$\pm$0.45 s (2.38$\pm$0.45 s) 2.6$\pm$0.96 s (2.5$\pm$0.82 s) 2.64$\pm$1.45 s (2.6$\pm$1.29 s)
Max. Latency 4.5 s (4.5 s) 6 s (5.25 s) 9.75 s (9.75 s)
@@
4.4 Threshold Effect in MAM
@@
Table 9: Effect of different activation thresholds on MAM
Threshold Sens Spec DFE Rejection Ratio Inference Time
@@
Figure 7: Impact of activity threshold on SSL performance
Fig. 7 shows how metrics like specificity and rejection ratio increase steadily with the threshold, while sensitivity and overall detection performance (F1 score) gradually decline. Careful selection of the threshold is crucial to balance the requirements of accurately detecting active periods while effectively rejecting inactive intervals to optimize computational resources.
Figure 8: Performance analysis for different amounts of labeled data.
Fig. 9 shows that as the activity threshold increases, the model execution time decreases for both supervised and self-supervised models. At a threshold of 1.2g, the SSL model’s execution time reaches approximately 1.1 ms (Table 9, Fig. 9).
Figure 9: Effect of activity threshold on inference time
This reduction is due to the MAM effectively filtering out inactive periods, reducing the need for the computationally intensive SSL model. The SSL model has slightly higher inference times than the supervised model due to an additional MLP layer on top of the pre-trained encoder. Despite the modest increase, the SSL model demonstrates enhanced robustness and superior performance using only 40-60% of labeled training data. The average execution time per window is 0.0295 ms for the supervised model and 0.0379 ms for the SSL model.
4.5 Effect of Labels During Pre-training
@@ 4.6 Post Processing
5 Discussion & Conclusion
In this study, we introduced LIFT-PD, a computationally simple self-supervised learning framework for real-time and patient-independent detection of Freezing of Gait (FoG) episodes in patients with Parkinson’s Disease (PD). LIFT-PD addresses critical challenges in continuous, in-home monitoring by eliminating the need for extensive labeled data, employing a unique differential hopping windowing technique to handle imbalanced datasets, and optimizing power consumption through opportunistic model activation. Our results show that LIFT-PD achieves performance comparable to the state-of-the-art methods while consuming less power and needing fewer annotated training samples, making it effective, generalizable, and suitable for deployment on wearable devices. Its ability to function with minimal data annotation and handle imbalanced datasets reduces the burden on healthcare providers, improving cost-effectiveness. It uses a single triaxial accelerometer, operates directly on raw sensor data without extensive preprocessing, and selectively activates the model only during active periods, significantly enhancing energy efficiency. By relying on a single, easily wearable sensor, the system enhances patient comfort, making it feasible for continuous, real-world monitoring in home environments.
From a clinical perspective, LIFT-PD can play a crucial role in the management of FoG, which is often unresponsive to traditional PD medications. Accurate detection of FoG events, as demonstrated by LIFT-PD, allows for the timely delivery of external cues, such as rhythmic tones or vibrotactile stimulation, which can help break or prevent FoG episodes without causing habituation. This targeted cueing approach enhances treatment efficacy and supports more personalized interventions [30, 3]. The system’s real-time, continuous monitoring capability using a single triaxial accelerometer makes it practical for in-home use, providing clinicians with more comprehensive data on the patient’s condition outside clinical settings.
Overall, LIFT-PD provides a practical, real-world solution for FoG detection, improving PD management, and improving patient quality of life. By providing continuous, real-time monitoring in home environments, the system supports more accurate assessments of a patient’s condition outside of clinical settings.

### v2

@@ 4.4. Compare with State-of-the-Art Models
One Class Classifier (naghavi2021towards) 90.8% 71.6% 0.86 0.77 0.89
Semi-Supervised Model (mikos2017real) 72.3% 0.71
Multi-head CNN (10.1016/j.artmed.2022.102459) 97.3% 94.6% 0.56 0.68 0.49
@@
4.5. Duration Based FoG Episode
@@
[…]
Table 10. Duration-based FoG episode analysis.
Duration
FoG Episodes & windows detection rate (%) with
latency (s) and standard deviation
Small, 0–6 s
Medium, 6–12 s
Large, >12 s
FoG Episodes
83.3%
(82.8%)
100%
(98.8%)
100%
(100%)
FoG Windows
68%
(71.4%)
81.9%
(84.1%)
91.1%
(92.6%)
Avg. Latency $\pm$ SD
[…]
Max. Latency
4.5 s
(4.5 s)
6 s
(5.25 s)
9.75 s
(9.75 s)
@@
Failure analysis for short FoG episodes: Short episodes (¡ 6 s) show a lower episode-level detection rate (83.3%, Table 10) mainly because their onsets and offsets often coincide with window boundaries. This reduces the proportion of FoG-dominant samples within a 3-s window, leading to ambiguous temporal cues and occasional delays. In addition, when the MAM threshold remains high at the start of a movement bout, model activation may be deferred, slightly delaying the first positive prediction. Future work will explore adaptive windowing (e.g., 2.5-s windows) and boundary-aware smoothing to improve responsiveness for brief episodes.
4.6. Threshold Effect in Model Activation Module (MAM)
@@
Table 11. Effect of different activation thresholds on MAM
Threshold
Sens Spec DFE
Rejection Ratio
Inference Time
@@
Figure 8. Impact of activity threshold on SSL performance
Figure 9. Effect of activity threshold on inference time
Fig. 8 shows how metrics like specificity and rejection ratio increase steadily with the threshold, while sensitivity and overall detection performance (F1 score) gradually decline. Careful selection of the threshold is crucial to balance the requirements of accurately detecting active periods while effectively rejecting inactive intervals to optimize computational resources. Fig. 9 shows that as the activity threshold increases, the model execution time decreases for both supervised and self-supervised models. At a threshold of 1.2g, the SSL model’s execution time reaches approximately 1.1 ms (Table 11, Fig. 9). This reduction is due to the MAM effectively filtering out inactive periods, reducing the need for the computationally intensive SSL model. The SSL model has slightly higher inference times than the supervised model due to an additional MLP layer on top of the pre-trained encoder. Despite the modest increase, the SSL model demonstrates enhanced robustness and superior performance using only 40-60% of labeled training data. The average execution time per window is 0.0295 ms for the supervised model and 0.0379 ms for the SSL model.
4.6.1. Battery-Life Estimation
Battery-life estimation was derived from measured inference times obtained on the local hardware used for experimentation. Each activity-threshold setting ($\alpha$) yielded a distinct average inference time (Table 11), which was used as a proxy for energy consumption. Assuming energy consumption scales linearly with inference duration, the average compute energy can be modeled as
$E_{\text{avg}}=(1-\rho)E_{\text{inf}}+E_{\text{base}},$
where $\rho$ denotes the rejection ratio and $E_{\text{inf}}$ corresponds to the measured inference energy (or time). At $\alpha=1.2\,g$, the rejection ratio reached $0.59$, reducing the inference time from $3.31\,\text{ms}$ to $1.1\,\text{ms}$—approximately a $2.4\times$ decrease in computational duty. This improvement translates to an estimated runtime extension from about 20 hours to over 48 hours on a comparable low-power wearable device. Because MAM operates on the mean-removed acceleration magnitude, it remains robust to sensor orientation, minor displacement, and brief motion artifacts.
4.7. Effect of Window Size in Episode Detection
Figure 10. Impact of window size on FoG detection latency across different episode durations (Small, Medium, Large).
We conducted extensive validation by randomly taking a 10% subset of the training data as a validation set. Our results indicated that longer windows, such as 4s or 5s, showed a delay in detecting FoG episodes, possibly due to including unnecessary gait segments not closely linked to FoG onset, in line with previous studies (mazilu2012online; koltermann2024gait). Among evaluated window lengths (2-5 seconds), the 3-second window was optimal, balancing detection latency and data relevance. Although intuitively shorter windows (2s) could reduce latency in some cases for larger FoG episodes, our validation revealed that shorter windows contained insufficient contextual information for reliable detection, thus slightly increasing the average detection latency compared to the 3-second window shown in Fig. 10. Longer windows (3.5s, 4s, 5s) further increased latency without accuracy gains. Therefore, a 3-second window was selected for optimal performance.
Figure 11. Performance analysis for different amounts of labeled data.
4.8. Effect of Labels During Pre-training
@@ 4.9. Post Processing
5. Limitations
While LIFT-PD demonstrates promising results and practical advantages for real-world deployment, several limitations remain. First, although the model generalizes well across diverse datasets and patient groups, further validation in larger patient cohorts with varying clinical conditions and disease progression stages, including patients exhibiting tremor-dominant or other atypical Parkinsonian motor symptoms, is needed to confirm robustness and clinical applicability. Specifically, additional assessment is required to determine the model’s accuracy in detecting FoG episodes in the presence of confounding symptoms such as resting tremors, dyskinesia, or bradykinesia, which often coexist in individuals with Parkinson’s disease. Second, while the system operates effectively with a single accelerometer, incorporating additional sensor modalities such as gyroscopes or physiological sensors (i.e., heart rate, skin conductance) might enhance detection accuracy, particularly for subtle or early-stage FoG events. Third, the Opportunistic Inference Module (OIM), though significantly improving energy efficiency, might occasionally delay activation, potentially missing very brief or subtle FoG events. Fourth, although decision thresholds in OIM were empirically optimized using a small validation subset from the training data, this threshold selection process remains dataset-specific and may not generalize across varying conditions or sensor configurations. We acknowledge that free-living environments introduce high variability, and are currently investigating dataset-agnostic calibration techniques using an independent dataset under acquisition. This future work aims to improve the robustness of threshold selection and ensure consistent performance across diverse deployment settings.
Finally, despite using significantly fewer labeled samples through self-supervised learning, the initial tuning of parameters and training still relies partially on labeled data, and completely unsupervised adaptation remains an open challenge. Future work should address these aspects to further optimize performance, clinical relevance, and adaptability of LIFT-PD.
6. Discussion
In this study, we introduced LIFT-PD, a computationally efficient and robust self-supervised learning framework designed specifically for real-time, patient-independent detection of Freezing of Gait (FoG) episodes in patients with Parkinson’s Disease (PD). Our proposed approach addresses key challenges inherent in continuous in-home FoG monitoring, particularly the reliance on large, extensively labeled datasets and substantial power consumption.
By implementing a novel Differential Hopping Windowing Technique (DHWT), LIFT-PD effectively handles imbalanced data and diverse gait variations, thereby reducing annotation burdens and enhancing clinical scalability. Furthermore, the integration of an opportunistic inference module substantially decreases power consumption by activating the deep learning model exclusively during ambulatory periods. Such selective activation translates to a significant improvement in battery life, making continuous (>>48-hour) wearable monitoring practical.
This system requires minimal data preprocessing, utilizing only a single triaxial accelerometer placed comfortably at the waist—an approach traditionally considered suboptimal yet proven highly effective through our self-supervised learning strategy. Clinically, accurate and real-time FoG detection as provided by LIFT-PD enables timely cue delivery (e.g., rhythmic auditory, visual, or vibrotactile stimuli), critical for breaking or preventing episodes without causing habituation (nieuwboer2008cueing; cosentino2023one).
To evaluate the clinical robustness of our model, we stratified performance by disease severity, age, gender, and medication state. These subgroups reflect known sources of variability in Parkinson’s Disease presentation and management. LIFT-PD consistently achieved high F1 scores and specificity across all groups, with particularly strong performance in older adults and those in the On-medication state. The system’s ability to robustly detect FoG episodes across varying patient demographics—including different severities, medication states, and age groups—further emphasizes its practical applicability for real-world clinical scenarios, enabling more personalized patient management.
Most mis-detections occur in short, transient FoG episodes (¡6s) where the transition dynamics dominate the window and movement magnitude is low. Such cases challenge both temporal continuity and MAM activation sensitivity. Adaptive thresholding and shorter, overlapping inference windows are promising directions to further reduce these errors. While the detection rate for short episodes is lower than for longer episodes, it is important to contextualize this performance within clinical relevance. Short FoG episodes (¡6 seconds) typically have less severe impact on patient mobility and fall risk compared to prolonged episodes. However, they may serve as early indicators of medication wearing off or precursors to longer episodes. The average detection latency of 2.42 ± 0.45 seconds for detected short episodes (Table 10) still provides a useful window for preventive cueing interventions.
In addition, we compared LIFT-PD against multiple state-of-the-art semi-supervised learning (Semi-SL) approaches, including the published Mikos et al. model (mikos2017real), a $\Pi$-model baseline, and a Mean-Teacher (MT) variant, all evaluated using the same data splits and architecture. While these Semi-SL methods leveraged unlabeled data through consistency regularization, they required a larger proportion of labeled samples to achieve stable performance and showed reduced sensitivity in detecting brief FoG events. In contrast, the proposed self-supervised learning (SSL) framework achieved superior recall and F1-scores while using fewer labeled data ( 40%), demonstrating its ability to learn robust, label-agnostic gait representations that generalize effectively across subjects and clinical conditions.
7. Conlusion
In conclusion, LIFT-PD presents a practical and clinically meaningful solution for real-time detection of Freezing of Gait episodes in Parkinson’s Disease. The system successfully achieves robust, patient-independent monitoring through innovative self-supervised learning and opportunistic inference strategies, reducing dependency on labeled data and significantly improving battery life. Its reliance on a single waist-worn accelerometer ensures patient comfort and adherence, thus enhancing feasibility for continuous at-home use. Furthermore, LIFT-PD demonstrates robust generalization across age, gender, disease severity, and medication states, and its outputs show significant correlation with clinical assessments (e.g., NFOGQ and UPDRS-III Off scores), reinforcing its translational potential. Ultimately, by delivering accurate and timely FoG detection and supporting targeted, personalized cue-based interventions, LIFT-PD significantly advances in-home monitoring capabilities, improves PD symptom management, and contributes positively to patients’ quality of life.
Data & Code Availability
This study utilized the publicly available tDCS FoG dataset. The dataset comprises wearable 3D lower-back sensor recordings annotated for freezing of gait (FoG) episodes in individuals with Parkinson’s disease. Ground truth labels for three event classes—Start Hesitation, Turn, and Walking—were derived through expert frame-by-frame video annotation. Access to the dataset is available via the competition website: https://kaggle.com/competitions/tlvmc-parkinsons-freezing-gait-prediction
The codebase for the LIFT-PD framework, including model implementation, training scripts, and evaluation pipelines, is publicly available at: https://github.com/shovito66/LIFT-PD